A photon has an energy of 1.58x10-18 Joules. What is the wavelength (in nm) of this light?
1
Expert's answer
2019-10-04T04:37:14-0400

E = hv= hc/lambda

Lambda= hc/E= 6.62e-34 x 3e8/1.58e-18=1.25696203e-7 m.

1.25696203e-7m= 125.696203 nm.